# Beam Enumeration: Probabilistic Explainability For Sample Efficient Self-conditioned Molecular Design

### v1

@@ C.2 Extracted Substructures
Figure C11: Substructures extracted in the illustrative example at varying epochs based on Structure and Scaffold.
@@
To illustrate the capability of Beam Enumeration to extract meaningful substructures, Fig. C11 shows the top 5 most probable substructures at epochs 1, 20, 40, 60, 80, and 100 based on Structure (extract any valid structure) and Scaffold (extract valid Bemis-MurckoBemis and Murcko (1996) scaffold) using a top $k$ of 2 and 18 beam steps. We make two crucial observations here. Firstly, Structure often extracts small functional groups which makes the self-conditioned filtering much more permissive as it is more likely for a molecule to possess a specific functional group than a specific scaffold. Secondly, benzene appears often and perhaps unsurprisingly as it is ubiquitous in nature. Based on these observations, we design Beam Enumeration to only extract substructures containing at least one heteroatom on the assumption that heteroatoms are much more informative in forming polar interactions in drug molecules, e.g., a hydrogen-bond cannot form from benzene. Finally, the general observation is that the most probable substructures gradually contain more heteroatoms, as desired.
C.3 Supplementary Main Text Results
@@
In this section, we present the same table as the main text illustrative experiment. The only difference is that the IntDiv1Polykovskiy et al. (2020b) is also annotated in the table here to show that the sample efficiency improvements of Beam Enumeration come only at a small trade-off in diversity (Table 4). In agreement with our observations in the hyperparameters grid search (Appendix A), Structure extraction with ’Structure Minimum Size’ enforcement leads to highly specific substructures which decrease diversity relative to Scaffold extraction but with potential gains in sample efficiency as evidenced in the drug discovery case studies (Appendix D). We further perform statistical testing using Welch’s t-test to compare all metrics for Scaffold with ’Structure Minimum Size’ = 15 and Baseline Augmented MemoryGuo and Schwaller (2023). For the experiments that had unsuccessful replicates, we use the total number of successful experiments, e.g., Oracle Burden>10 (100), the Baseline was unsuccessful in 69/100 replicates so a 31 sample size was used. Overall, all p-values are significant at the 95% confidence level.
@@
Table 4: Illustrative experiment: Beam Enumeration improves the sample efficiency of Augmented Memory. All experiments were run for 100 replicates with an oracle budget of 5,000 calls and reported values are the mean and standard deviation. Scaffold and Structure indicate the type of substructure and the number after is the ’Structure Minimum Size’. Parentheses after Oracle Burden denote the cut-off number of molecules. Parentheses after values represent the number of unsuccessful replicates (for achieving the metric). The IntDiv1Polykovskiy et al. (2020a) is annotated under each Generative Yield. Welch’s t-test is used to compare the difference between Scaffold with ’Structure Minimum Size’ = 15 and Baseline Augmented MemoryGuo and Schwaller (2023). All p-values are significant.
@@
C.4 Beam Enumeration works in Exploitation Scenarios
Table 5: Beam Enumeration works in exploitation scenarios. all experiments were run for 100 replicates with an oracle budget of 5,000 calls and reported values are the mean and standard deviation. Parentheses after Oracle Burden denote the cut-off number of molecules. The IntDiv1Polykovskiy et al. (2020a) is annotated under each Generative Yield. Welch’s t-test is used to compare the difference between Scaffold with ’Structure Minimum Size’ = 15 and Baseline Augmented MemoryGuo and Schwaller (2023). All p-values are significant.
@@ C.4 Beam Enumeration works in Exploitation Scenarios
In the main text illustrative experiment, Augmented MemoryGuo and Schwaller (2023) was used with Selective Memory Purge activated which is the mechanism to promote chemical space exploration, as described in the original work. For completeness, we show that Beam Enumeration also works in pure exploitation scenarios where the goal is only to generate high reward molecules even if the same molecule is repeatedly sampled (Table 5). We perform statistical testing using Welch’s t-test to compare all metrics for Scaffold with ’Structure Minimum Size’ = 15 and Baseline Augmented MemoryGuo and Schwaller (2023). For the experiments that had unsuccessful replicates, we use the total number of successful experiments, e.g., Oracle Burden>10 (100), the Baseline was unsuccessful in 69/100 replicates so a 31 sample size was used. Overall, all p-values are significant at the 95% confidence level.
C.5 Self-conditioned Filtering: Structure vs Scaffold
@@
Figure C12: Behaviour of Beam Enumeration using Structure and Scaffold on self-conditioning.
@@
There is a clear discrepancy in the substructures extracted by Structure and Scaffold. In particular, Structure substructures contain small functional groups which is much more permissive when used as a filter criterion compared to full scaffolds. Therefore, one would expect that many molecules in the sampled batches would be kept when using Structure Beam Enumeration. We plot the average number of molecules kept out of 64 (batch size) across the generative run when using Beam Enumeration. Note that the experiments ran for variable epochs due to the stochasticity of Beam Enumeration self-filtering. The number of epochs shown in C12 is the minimum number of epochs out of 100 replicates. Therefore, the average values shown are averaged over 100 replicates. It is evident that Structure is more lenient as many generated molecules make it through the filter compared to Scaffold which maintains a relatively strict filter. One interesting observation is that self-conditioning does not lead to obvious mode collapse. Self-conditioning is inherently biased and one would be concerned that the model gets stuck at generating the same molecules repeatedly. The fact that self-conditioning with Scaffold continues to filter throughout the entire generative run shows that the model is continually moving to new chemical space, supporting findings from the original Augmented MemoryGuo and Schwaller (2023) work that Selective Memory Purge (built-in diversity mechanism) is capable of preventing mode collapse.
Appendix D Drug Discovery Case Studies
@@
This section contains information on the Autodock VinaTrott and Olson (2010) docking protocol from receptor grid preparation to docking execution. The Beam Enumeration hyperparameters grid search results are presented for all three drug discovery case studies followed by analysis. Examples of extracted substructures are also shown and commentary provided to their significance and explainability. Finally, the wall times of all experiments are presented.
D.1 AutoDock Vina Receptor Preparation and Docking
@@
All docking grids were prepared using DockStreamGuo et al. (2021b) which uses PDBFixerEastman et al. (2017) to refine receptor structures. The search box for all grids was 15Å x 15Å x 15Å. Docking was also performed through DockStream and followed a two step process: conformer generation using the RDKit Universal Force Field (UFF)Rappé et al. (1992) with the maximum convergence set to 600 iterations and then Vina docking was parallelized over 36 CPU cores (Intel(R) Xeon(R) Platinum 8360Y processors).
DRD2 - Dopamine Type 2 Receptor. The PDB ID is 6CM4Wang et al. (2018) and the docking grid was centered at (x, y, z) = (9.93, 5.85, -9.58).
MK2 - MK2 Kinase. The PDB ID is 3KC3Argiriadi et al. (2010) and one monomer was extracted. The docking grid for the extracted monomer was centered at (x, y, z) = (-61.62, 30.31, -21.9).
AChE - Acetylcholinesterase. The PDB ID is 1EVEKryger et al. (1999) and the docking grid was centered at (x, y, z) = (2.78, 64.38, 67.97).
D.2 Beam Enumeration Hyperparameters Grid Search Results
@@
  • Optimization Algorithm = [Augmented MemoryGuo and Schwaller (2023), REINVENTOlivecrona et al. (2017); Blaschke et al. (2020b)]
@@
  • Token Sampling Method = ["topk", "sample"]
All hyperparameter combinations (8) were tried and run for 3 replicates each for statistical reproducibility, total of 144 experiments. There are two main results we want to convey: firstly, the optimal hyperparameters are the same for all three drug discovery case studies and only the Substructure Type differs between the optimal hyperparameters here and the illustrative experiment. Secondly, Beam Enumeration is a task-agnostic general method that can be applied to existing algorithms including Augmented MemoryGuo and Schwaller (2023) and REINVENTOlivecrona et al. (2017); Blaschke et al. (2020a). At the end of this section, we present these hyperparameters and designate these the default values. All grid search results are now presented in following tables:
Table 6: DRD2Wang et al. (2018) case study hyperparameters grid search results for Augmented MemoryGuo and Schwaller (2023). All experiments were run in triplicate and the reported values are the mean and standard deviation. "Sample" denotes "sample" token sampling. All metrics are for the reward threshold > 0.8. The IntDiv1Polykovskiy et al. (2020a) is annotated under Generative Yield. * and ** denote one and two replicates were unsuccessful, respectively.
@@
Table 7: DRD2Wang et al. (2018) case study hyperparameters grid search results for REINVENTOlivecrona et al. (2017); Blaschke et al. (2020a). All experiments were run in triplicate and the reported values are the mean and standard deviation. "Sample" denotes "sample" token sampling. The IntDiv1Polykovskiy et al. (2020a) is annotated under Generative Yield. All metrics are for the reward threshold > 0.8. * and ** denote one and two replicates were unsuccessful, respectively.
@@
Table 8: MK2Argiriadi et al. (2010) case study hyperparameters grid search results for Augmented MemoryGuo and Schwaller (2023). All experiments were run in triplicate and the reported values are the mean and standard deviation. "Sample" denotes "sample" token sampling. All metrics are for the reward threshold > 0.8. The IntDiv1Polykovskiy et al. (2020a) is annotated under Generative Yield. * and ** denote one and two replicates were unsuccessful, respectively.
@@
Table 9: MK2Argiriadi et al. (2010) case study hyperparameters grid search results for REINVENTOlivecrona et al. (2017); Blaschke et al. (2020a). All experiments were run in triplicate and the reported values are the mean and standard deviation. "Sample" denotes "sample" token sampling. All metrics are for the reward threshold > 0.8. The IntDiv1Polykovskiy et al. (2020a) is annotated under Generative Yield. * and ** denote one and two replicates were unsuccessful, respectively.
@@
Table 10: AChEKryger et al. (1999) case study hyperparameters grid search results for Augmented memoryGuo and Schwaller (2023). All experiments were run in triplicate and the reported values are the mean and standard deviation. "Sample" denotes "sample" token sampling. All metrics are for the reward threshold > 0.8. The IntDiv1Polykovskiy et al. (2020a) is annotated under Generative Yield. * and ** denote one and two replicates were unsuccessful, respectively.
@@
Table 11: AChEKryger et al. (1999) case study hyperparameters grid search results for REINVENTOlivecrona et al. (2017); Blaschke et al. (2020a). All experiments were run in triplicate and the reported values are the mean and standard deviation. "Sample" denotes "sample" token sampling. The IntDiv1Polykovskiy et al. (2020a) is annotated under Generative Yield. All metrics are for the reward threshold > 0.8. * and ** denote one and two replicates were unsuccessful, respectively.
@@
Structure Sample Size 15 316 $\pm$ 253 190 $\pm$ 146 125 $\pm$ 50 561 $\pm$ 140 2683 $\pm$ 320
@@
Based on the results from the hyperparameters grid search in the drug discovery case studies, we make two key observations: firstly, Structure extraction with ’Structure Minimum Size’ = 15 is now the most performant, on average (for both Augmented MemoryGuo and Schwaller (2023) and REINVENTOlivecrona et al. (2017); Blaschke et al. (2020a)). This is in contrast to Scaffold extraction in the illustrative experiment which we rationalize through the permissive nature of the experiment compared to the docking experiments which require structure specificity. Previously, small deviations in the substructures may not have a significant impact on the reward. In physics-based oracles such as VinaTrott and Olson (2010) docking used here, small substructure differences can have an enormous impact on the outcome since the pose requires specific complementary to the protein binding site. The second observation we make which is in agreement with the illustrative experiment is that "sample" token sampling has more variance and does not perform better than "topk". The rationale is the same in that docking requires specificity and lower probability substructures exhibit more variable performance. Based on all the observations from the illustrative experiment and the drug discovery case studies, we designate the following default hyperparameter values:
@@
  • Token Sampling Method = "topk"
D.3 Examples of Extracted Substructures: Structure Extraction with ’Structure Minimum Size’ = 15
In this section, the top substructures at the end of the generative experiments (using Augmented MemoryGuo and Schwaller (2023)) are shown for all three drug discovery case studies (3 replicates). All experiments are for Structure extraction with ’Structure Minimum Size’ = 15. The extracted substructures are commonly scaffolds with "branch points", i.e., a central scaffold with single carbon bond extensions outward, which heavily bias generation. We posit that this may be a reason why Structure extraction can be more performant than Scaffold, as observed in the hyperparameters grid search in the previous subsection.
Figure D13: Augmented MemoryGuo and Schwaller (2023) DRD2Wang et al. (2018) substructures with Structure extraction and ’Structure Minimum Size’ = 15 after 5,000 oracle calls.
Figure D14: Augmented MemoryGuo and Schwaller (2023) MK2Argiriadi et al. (2010) substructures with Structure extraction and ’Structure Minimum Size’ = 15 after 5,000 oracle calls.
Figure D15: Augmented MemoryGuo and Schwaller (2023) AChEKryger et al. (1999) substructures with Structure extraction and ’Structure Minimum Size’ = 15 after 5,000 oracle calls.
D.4 Wall Times
Table 12: Wall times for all drug discovery case studies hyperparameters grid search using Augmented MemoryGuo and Schwaller (2023) and REINVENTOlivecrona et al. (2017); Blaschke et al. (2020a). "Sample" denotes "sample" token sampling. All experiments were run in triplicate and the values are the mean and standard deviation.

### v2

@@ D.2 Extracted Substructures
Figure D11: Substructures extracted in the illustrative example at varying epochs based on Structure and Scaffold.
@@
To illustrate the capability of Beam Enumeration to extract meaningful substructures, Fig. D11 shows the top 5 most probable substructures at epochs 1, 20, 40, 60, 80, and 100 based on Structure (extract any valid structure) and Scaffold (extract valid Bemis-Murcko (Bemis & Murcko (1996) scaffold) using a top $k$ of 2 and 18 beam steps. We make two crucial observations here. Firstly, Structure often extracts small functional groups which makes the self-conditioned filtering much more permissive as it is more likely for a molecule to possess a specific functional group than a specific scaffold. Secondly, benzene appears often and perhaps unsurprisingly as it is ubiquitous in nature. Based on these observations, we design Beam Enumeration to only extract substructures containing at least one heteroatom on the assumption that heteroatoms are much more informative in forming polar interactions in drug molecules, e.g., a hydrogen-bond cannot form from benzene. Finally, the general observation is that the most probable substructures gradually contain more heteroatoms, as desired.
D.3 Supplementary Main Text Results
@@
In this section, we present the same table as the main text illustrative experiment. The only difference is that the IntDiv1 (Polykovskiy et al. (2020a) is also annotated in the table here to show that the sample efficiency improvements of Beam Enumeration come only at a small trade-off in diversity (Table 4). In agreement with our observations in the hyperparameters grid search (Appendix B), Structure extraction with ’Structure Minimum Size’ enforcement leads to highly specific substructures which decrease diversity relative to Scaffold extraction but with potential gains in sample efficiency as evidenced in the drug discovery case studies (Appendix E). We further perform statistical testing using Welch’s t-test to compare all metrics for Scaffold with ’Structure Minimum Size’ = 15 and Baseline Augmented Memory (Guo & Schwaller (2023). For the experiments that had unsuccessful replicates, we use the total number of successful experiments, e.g., Oracle Burden>10 (100), the Baseline was unsuccessful in 69/100 replicates so a 31 sample size was used. Overall, all p-values are significant at the 95% confidence level.
@@
Table 4: Illustrative experiment: Beam Enumeration improves the sample efficiency of Augmented Memory. All experiments were run for 100 replicates with an oracle budget of 5,000 calls and reported values are the mean and standard deviation. Scaffold and Structure indicate the type of substructure and the number after is the ’Structure Minimum Size’. Parentheses after Oracle Burden denote the cut-off number of molecules. Parentheses after values represent the number of unsuccessful replicates (for achieving the metric). The IntDiv1 (Polykovskiy et al. (2020b) is annotated under each Generative Yield. Welch’s t-test is used to compare the difference between Scaffold with ’Structure Minimum Size’ = 15 and Baseline Augmented Memory (Guo & Schwaller (2023). All p-values are significant.
@@
D.4 Beam Enumeration works in Exploitation Scenarios
Table 5: Beam Enumeration works in exploitation scenarios. all experiments were run for 100 replicates with an oracle budget of 5,000 calls and reported values are the mean and standard deviation. Parentheses after Oracle Burden denote the cut-off number of molecules. The IntDiv1 (Polykovskiy et al. (2020b) is annotated under each Generative Yield. Welch’s t-test is used to compare the difference between Scaffold with ’Structure Minimum Size’ = 15 and Baseline Augmented Memory (Guo & Schwaller (2023). All p-values are significant.
@@ D.4 Beam Enumeration works in Exploitation Scenarios
In the main text illustrative experiment, Augmented Memory (Guo & Schwaller (2023) was used with Selective Memory Purge activated which is the mechanism to promote chemical space exploration, as described in the original work. For completeness, we show that Beam Enumeration also works in pure exploitation scenarios where the goal is only to generate high reward molecules even if the same molecule is repeatedly sampled (Table 5). We perform statistical testing using Welch’s t-test to compare all metrics for Scaffold with ’Structure Minimum Size’ = 15 and Baseline Augmented Memory (Guo & Schwaller (2023). For the experiments that had unsuccessful replicates, we use the total number of successful experiments, e.g., Oracle Burden>10 (100), the Baseline was unsuccessful in 69/100 replicates so a 31 sample size was used. Overall, all p-values are significant at the 95% confidence level.
D.5 Self-conditioned Filtering: Structure vs Scaffold
@@
Figure D12: Behaviour of Beam Enumeration using Structure and Scaffold on self-conditioning.
@@
There is a clear discrepancy in the substructures extracted by Structure and Scaffold. In particular, Structure substructures contain small functional groups which is much more permissive when used as a filter criterion compared to full scaffolds. Therefore, one would expect that many molecules in the sampled batches would be kept when using Structure Beam Enumeration. We plot the average number of molecules kept out of 64 (batch size) across the generative run when using Beam Enumeration. Note that the experiments ran for variable epochs due to the stochasticity of Beam Enumeration self-filtering. The number of epochs shown in D12 is the minimum number of epochs out of 100 replicates. Therefore, the average values shown are averaged over 100 replicates. It is evident that Structure is more lenient as many generated molecules make it through the filter compared to Scaffold which maintains a relatively strict filter. One interesting observation is that self-conditioning does not lead to obvious mode collapse. Self-conditioning is inherently biased and one would be concerned that the model gets stuck at generating the same molecules repeatedly. The fact that self-conditioning with Scaffold continues to filter throughout the entire generative run shows that the model is continually moving to new chemical space, supporting findings from the original Augmented Memory (Guo & Schwaller (2023) work that Selective Memory Purge (built-in diversity mechanism) is capable of preventing mode collapse.
Appendix E Drug Discovery Case Studies
@@
This section contains information on the Autodock Vina (Trott & Olson (2010) docking protocol from receptor grid preparation to docking execution. The Beam Enumeration hyperparameters grid search results are presented for all three drug discovery case studies followed by analysis. Examples of extracted substructures are also shown and commentary provided to their significance and explainability. Finally, the wall times of all experiments are presented.
E.1 AutoDock Vina Receptor Preparation and Docking
@@
All docking grids were prepared using DockStream (Guo et al. (2021a) which uses PDBFixer (Eastman et al. (2017) to refine receptor structures. The search box for all grids was 15Å x 15Å x 15Å. Docking was also performed through DockStream and followed a two step process: conformer generation using the RDKit Universal Force Field (UFF) (Rappé et al. (1992) with the maximum convergence set to 600 iterations and then Vina docking was parallelized over 36 CPU cores (Intel(R) Xeon(R) Platinum 8360Y processors).
DRD2 - Dopamine Type 2 Receptor. The PDB ID is 6CM4 (Wang et al. (2018) and the docking grid was centered at (x, y, z) = (9.93, 5.85, -9.58).
MK2 - MK2 Kinase. The PDB ID is 3KC3 (Argiriadi et al. (2010) and one monomer was extracted. The docking grid for the extracted monomer was centered at (x, y, z) = (-61.62, 30.31, -21.9).
AChE - Acetylcholinesterase. The PDB ID is 1EVE (Kryger et al. (1999) and the docking grid was centered at (x, y, z) = (2.78, 64.38, 67.97).
E.2 Beam Enumeration Hyperparameters Grid Search Results
@@
  • Optimization Algorithm = [Augmented Memory (Guo & Schwaller (2023), REINVENT (Olivecrona et al. (2017); Blaschke et al. (2020)]
@@
  • Token Sampling Method = [”topk”, ”sample”]
All hyperparameter combinations (8) were tried and run for 3 replicates each for statistical reproducibility, total of 144 experiments. There are two main results we want to convey: firstly, the optimal hyperparameters are the same for all three drug discovery case studies and only the Substructure Type differs between the optimal hyperparameters here and the illustrative experiment. Secondly, Beam Enumeration is a task-agnostic general method that can be applied to existing algorithms including Augmented Memory (Guo & Schwaller (2023) and REINVENT (Olivecrona et al. (2017); Blaschke et al. (2020). At the end of this section, we present these hyperparameters and designate these the default values. All grid search results are now presented in following tables:
Table 6: DRD2 (Wang et al. (2018) case study hyperparameters grid search results for Augmented Memory (Guo & Schwaller (2023). All experiments were run in triplicate and the reported values are the mean and standard deviation. ”Sample” denotes ”sample” token sampling. All metrics are for the reward threshold >0.8. The IntDiv1 (Polykovskiy et al. (2020b) is annotated under Generative Yield. * and ** denote one and two replicates were unsuccessful, respectively.
@@
Table 7: DRD2 (Wang et al. (2018) case study hyperparameters grid search results for REINVENT (Olivecrona et al. (2017); Blaschke et al. (2020). All experiments were run in triplicate and the reported values are the mean and standard deviation. ”Sample” denotes ”sample” token sampling. The IntDiv1 (Polykovskiy et al. (2020b) is annotated under Generative Yield. All metrics are for the reward threshold >0.8. * and ** denote one and two replicates were unsuccessful, respectively.
@@
Table 8: MK2 (Argiriadi et al. (2010) case study hyperparameters grid search results for Augmented Memory (Guo & Schwaller (2023). All experiments were run in triplicate and the reported values are the mean and standard deviation. ”Sample” denotes ”sample” token sampling. All metrics are for the reward threshold >0.8. The IntDiv1 (Polykovskiy et al. (2020b) is annotated under Generative Yield. * and ** denote one and two replicates were unsuccessful, respectively.
@@
Table 9: MK2 (Argiriadi et al. (2010) case study hyperparameters grid search results for REINVENT (Olivecrona et al. (2017); Blaschke et al. (2020). All experiments were run in triplicate and the reported values are the mean and standard deviation. ”Sample” denotes ”sample” token sampling. All metrics are for the reward threshold >0.8. The IntDiv1 (Polykovskiy et al. (2020b) is annotated under Generative Yield. * and ** denote one and two replicates were unsuccessful, respectively.
@@
Table 10: AChE (Kryger et al. (1999) case study hyperparameters grid search results for Augmented memory (Guo & Schwaller (2023). All experiments were run in triplicate and the reported values are the mean and standard deviation. ”Sample” denotes ”sample” token sampling. All metrics are for the reward threshold >0.8. The IntDiv1 (Polykovskiy et al. (2020b) is annotated under Generative Yield. * and ** denote one and two replicates were unsuccessful, respectively.
@@
Table 11: AChE (Kryger et al. (1999) case study hyperparameters grid search results for REINVENT (Olivecrona et al. (2017); Blaschke et al. (2020). All experiments were run in triplicate and the reported values are the mean and standard deviation. ”Sample” denotes ”sample” token sampling. The IntDiv1 (Polykovskiy et al. (2020b) is annotated under Generative Yield. All metrics are for the reward threshold >0.8. * and ** denote one and two replicates were unsuccessful, respectively.
@@
Structure Sample Size 15 316 $\pm$ 253 190 $\pm$ 146 125 $\pm$ 50 561 $\pm$ 140 2683 $\pm$ $*{{}^{*}}$
@@
Based on the results from the hyperparameters grid search in the drug discovery case studies, we make two key observations: firstly, Structure extraction with ’Structure Minimum Size’ = 15 is now the most performant, on average (for both Augmented Memory (Guo & Schwaller (2023) and REINVENT (Olivecrona et al. (2017); Blaschke et al. (2020)). This is in contrast to Scaffold extraction in the illustrative experiment which we rationalize through the permissive nature of the experiment compared to the docking experiments which require structure specificity. Previously, small deviations in the substructures may not have a significant impact on the reward. In physics-based oracles such as Vina (Trott & Olson (2010) docking used here, small substructure differences can have an enormous impact on the outcome since the pose requires specific complementary to the protein binding site. The second observation we make which is in agreement with the illustrative experiment is that ”sample” token sampling has more variance and does not perform better than ”topk”. The rationale is the same in that docking requires specificity and lower probability substructures exhibit more variable performance. Based on all the observations from the illustrative experiment and the drug discovery case studies, we designate the following default hyperparameter values:
@@
  • Token Sampling Method = ”topk”
E.3 Examples of Extracted Substructures: Structure Extraction with ’Structure Minimum Size’ = 15
In this section, the top substructures at the end of the generative experiments (using Augmented Memory (Guo & Schwaller (2023)) are shown for all three drug discovery case studies (3 replicates). All experiments are for Structure extraction with ’Structure Minimum Size’ = 15. The extracted substructures are commonly scaffolds with ”branch points”, i.e., a central scaffold with single carbon bond extensions outward, which heavily bias generation. We posit that this may be a reason why Structure extraction can be more performant than Scaffold, as observed in the hyperparameters grid search in the previous subsection.
Figure E13: Augmented Memory (Guo & Schwaller (2023) DRD2 (Wang et al. (2018) substructures with Structure extraction and ’Structure Minimum Size’ = 15 after 5,000 oracle calls.
Figure E14: Augmented Memory (Guo & Schwaller (2023) MK2 (Argiriadi et al. (2010) substructures with Structure extraction and ’Structure Minimum Size’ = 15 after 5,000 oracle calls.
Figure E15: Augmented Memory (Guo & Schwaller (2023) AChE (Kryger et al. (1999) substructures with Structure extraction and ’Structure Minimum Size’ = 15 after 5,000 oracle calls.
E.4 Wall Times
Table 12: Wall times for all drug discovery case studies hyperparameters grid search using Augmented Memory (Guo & Schwaller (2023) and REINVENT (Olivecrona et al. (2017); Blaschke et al. (2020). ”Sample” denotes ”sample” token sampling. All experiments were run in triplicate and the values are the mean and standard deviation.